OpenStudy (anonymous):

sin^2x - 2sinx + 1 = 0

OpenStudy (anonymous):

Solve for x??

OpenStudy (anonymous):

First solve for sin x, then go to the unit circle and find anywhere where sin equals x

OpenStudy (campbell_st):

this is an equation that is reducible to a quadratic. its actually a perfect square (sin(x) - 1)^2 = 0 now solve for x

OpenStudy (anonymous):

^^ if you wanna solve for sinx from the above answer \[(sinx-1)^{2}\]=0 then introduce a squareroot in both sides to get : \[\sin x-1=0\]

OpenStudy (anonymous):

Do you get what I am saying above?

OpenStudy (anonymous):

So would you then add 1 to both sides, get sinx = 1, then according to the unit circle the answer would be pi/2 ?

OpenStudy (anonymous):

it's for general solutions so actually pi/2 + 2pi n

OpenStudy (anonymous):

I think that so

OpenStudy (anonymous):

Yea you sinx will surely be equal to 1(one)

OpenStudy (anonymous):

so (sinx - 1)^2 = 0 when worked out equals sin x - 1?

OpenStudy (anonymous):

\[(sinx-1)^{2}=0 \] \[\sqrt{(sinx-1)^{2}}=\sqrt{0}\] \[(sinx-1)^{2*1/2} =0\] \[\sin x-1=0\] and therefore \[\sin x=1\]

OpenStudy (anonymous):

I got sinx^2 + 1 when I worked it out